OpenStudy (anonymous):

Nitrogen has a density of 1.17 x 10^-3 g/cm^3. If you have 2,500 cm^3 of nitrogen gas, which of the following is the correct mass of nitrogen, using significant figures? A. 2.92 g B. 2.2 x 10^6g C. 14 x 10^6g D. 2.9 g ***My answer is D. 2.9 g.... am i right? if not, please explain!! :) thx!

OpenStudy (anonymous):

Hi!! thanks for coming!!! :) sooo what do u think? :D

OpenStudy (dumbcow):

well to get mass you do: mass = density*volume mass = (1.17*10^-3) * 2500 = 2.925

OpenStudy (dumbcow):

hold on, let me figure out the sig figs :)

OpenStudy (anonymous):

oh and can u please double check another question after this one??? Here it is: An enzyme works best at 98.6°F. But the equation that you must use to describe it requires the temperature to be in K. Which of the following is the correct temperature? ([°F x 0.555] + 255.37 = K and [K -255.37] x 1.8 = °F) A. 37 K B. 482 K C. 310 K D. 236 K ***I think the answer is C. 310 K.... Am i right?? :)

OpenStudy (anonymous):

okay sure thing :)

OpenStudy (anonymous):

oh and if it were to be 2.925, wouldn't that round up to 2.93?? And that is not an answer choice right? :)

OpenStudy (dumbcow):

yes, however since 2500 only has 2 sig figs, the answer must only have 2 sig figs -->2.9

OpenStudy (anonymous):

YAy! so for the first problem, i was right?!? the answer was D. 2.9 g??? And also, is my second problem right?? i posted it above :)

OpenStudy (dumbcow):

yes you are right, its 310K

OpenStudy (anonymous):

YAy!! :) THANks for the help @dumbcow!! :) BTW i like ur username :)

OpenStudy (dumbcow):

yw and thanks